# QCS:Feature Refining from Quadruplet Cross Similarity for Facial Expression Recognition

### v1

@@ Introduction
Table 3 compares our best results with the current state-of-the-art methods on RAF-DB, AffectNet-7, and FERPlus. APViT and the other methods listed below are Transformer-based methods, while POSTER++, LA-Net, and S2D incorporate additional landmark information. Our proposed method achieves these results without utilizing extra landmark information or other extra training data, ensuring no inference time increase by removing all interactive modules and retaining a single branch. While QCS effectively separates redundant features, the heightened complexity of the interactive module may pose challenges during training, particularly for AffectNet-7 where over-sampling of underrepresented classes can exacerbate overfitting, ultimately leading to a performance trade-off. Nevertheless, our method obtains competitive results on AffectNet-7 and FERPlus when compared to Landmark-Aware methods, outperforming recent state-of-the-art methods on RAF-DB.
Methods RAF-DB FERPlus AffectNet-7
SCN(2020) 87.03 88.01 60.23
RUL(2021) 88.98 - -
DMUE(2021) 89.42 89.51 63.11
TransFER(2021) 90.91 90.83 65.20
Face2Exp(2022) 88.54 - 64.23
EAC(2022) 89.99 89.64 65.32
APViT(2022) 91.98 90.86 66.91
TAN(2023) 90.87 91.00 66.45
POSTER(2023) 92.05 91.62 67.31
POSTER++(2023) 92.21 - 67.49
S2D(2024) 92.21 91.01 66.42
DCS(Ours) 92.57 91.25 67.29
QCS(Ours) 92.47 91.21 66.91
QCS(Ours)‡ 92.83 91.50 -
Table 3: Performance comparison (%) on RAF-DB, AffectNet-7, and FERPlus. ‡means pre-trained on the AffectNet-8(Acc. 63.9).
@@
To visually demonstrate the effectiveness of our method in extracting distinctive features and separating redundant features, we calculated the cross similarity between every pair of the 7 expressions. We utilized the QCS described in Section 3 to simultaneously compute the similarity matrices between the anchor and positive as well as between the anchor and negative. Fig. 7 only displays the similarity heatmap used to calculate the anchor attention, and the heatmaps for positive and negative attention are not shown.
The heatmaps in diagonal positions show that the distinctive features extracted from images of the same category mainly concentrate on the areas of eyes and mouth corners. For example, Happy is mainly reflected in the areas of mouth corners and cheeks, while Disgust mainly in the areas of mouth and nasolabial folds. The exaggerated facial expressions of Anger are reflected in the entire facial features area, and the similar enlargement of eyes and mouths in Surprise and Fear results in similar heatmaps. The heatmaps in non-diagonal positions show the redundant features separated from comparing the images in a certain column (Key) with various expressions. As can be seen from the last row of heatmaps, Neutral, which does not have the facial features of expressions, mainly separates out feature regions from various expressions in the corners of the images.
Figure 7: The confusion matrix of Cross Similarity heatmaps on RAF-DB. The anchor image located above a certain column is used as the key to calculate cross similarity on that anchor. The yellow highlighted area represents high similarity.
@@ CONCLUSION
This paper proposes a feature refining network named Quadruplet Cross Similarity(QCS) for FER. The QCS comprises a closed-loop symmetric network formed by four weight-sharing branches and four cross modules for stable joint training. It extracts discriminative features within the same class and synchronously separate interfering features across different classes. The cross module is based on Cross Similarity Attention (CSA), a novel global spatial attention designed for different image pairs, which is input-output position-sensitive. Extensive experiments on several public datasets indicate that our QCS effectively refines the features into a purer form and achieves state-of-the-art performance. Furthermore, effectively transferring the cross-module’s performance back to the basic inference network while avoiding overfitting and ensuring the stability of the model’s performance remains a crucial task.
@@ Acknowledgments
This work was supported by funding from the National Natural Science Foundation of China (U21B2035 and 82372218).

### v2

@@ Introduction
Taking into account the potential overfitting issue in the cross module, we incorporated data augmentation such as random cropping, rotation and color jitter. We also optimized the sampling strategy for positive and negative samples, resulting in further improvements in model performance. (See the Appendix for details.)
Table 3 compares our best results with the current state-of-the-art methods on RAF-DB, FERPlus, and AffectNet-7/8. APViT and the other methods listed below are Transformer-based methods, while POSTER, POSTER++, and S2D incorporate additional landmark information. Our proposed method achieves these results solely by leveraging self-supervised contrastive learning to mine richer intrinsic information, without incorporating any extra landmark information.
While QCS effectively separates redundant features, its performance does not demonstrate a significant advantage compared to DCS. This may be due to the fact that more complex interactive modules may pose challenges during training, where over-sampling of underrepresented classes can exacerbate overfitting, ultimately resulting in a performance trade-off. Nevertheless, our method outperforms recent state-of-the-art methods on those datasets.
Methods RAF-DB FERPlus AffectNet-7 AffectNet-8
SCN(2020) 87.03 88.01 - 60.23
RUL(2021) 88.98 - - -
DMUE(2021) 89.42 89.51 - 63.11
TransFER(2021) 90.91 90.83 66.23 -
Face2Exp(2022) 88.54 - 64.23 -
EAC(2022) 89.99 89.64 65.32 -
APViT(2022) 91.98 90.86 66.91 -
TAN(2023) 90.87 91.00 66.45 -
POSTER(2023) 92.05 91.62 67.31 63.34
POSTER++(2023) 92.21 - 67.49 63.77
S2D‡(2024) 92.57 91.17 67.62 63.06
DCS(Ours) 92.57 91.41 67.66 64.40
QCS(Ours) 92.50 91.41 67.94 64.30
QCS(Ours)‡ 93.02 91.85 - -
Table 3: Performance comparison (%) on RAF-DB, AffectNet-7, and FERPlus. ‡means pre-trained on the AffectNet-8.
@@
Cross Similarity Visualization
To visually demonstrate the effectiveness of QCS in extracting distinctive features and separating redundant features, we calculated the cross similarity between every pair of the 7 expressions. We computed the similarity matrices between the anchor and positive as well as between the anchor and negative. Fig. 8 only displays the similarity heatmaps of the anchor that is used as Key, and the heatmaps for positive and negative are not shown.
Figure 8: The confusion matrix of Cross Similarity heatmaps on RAF-DB. The anchor image located above a certain column is used as the key to calculate cross similarity on that anchor. The yellow highlighted area represents high similarity. During adaptive fusion through Eq. 9, high-similarity features on the confusion matrix diagonal are enhanced, while those outside the diagonal are suppressed.
The heatmaps in diagonal positions show that the distinctive features extracted from images of the same category mainly concentrate on the areas of eyes and mouth corners. The heatmaps in non-diagonal positions show the redundant features separated from comparing the images in a certain column (Key) with various expressions. As can be seen from the last row of heatmaps, Neutral, which does not have the facial features of expressions, mainly separates out feature regions from various expressions in the corners of the images.
Figure 9: Cross Attentions visualization on RAF-DB. The QCS uniformly employs the Neutral image as the input of negative branch.
Attention Visualization
We conducted a visual comparison of the cross attentions in the Ablation Analysis. Due to the removal of the class token for cross module, we present SDPA by averaging the attention weights from all other tokens.From Row 2 of Fig. 9, it can be observed that the attention areas of SDPA are not prominent, and a significant number of redundant features in the corner regions of the images receive greater emphasis during information transmission. In contrast, CSA directly acts between the input and output feature vector of its own branch, enabling more direct feedback. As clearly shown in Row 3 of Fig. 9, our method achieves effective attention capture, with the effects mainly concentrated on key areas such as the facial features, cheeks, and eyebrows.
Rows 3 and 4 of Fig. 9 demonstrate the attentions of the cross module S in the DCS Network and the cross module SD in the QCS Network, respectively. It can be observed that the highlighted areas in QCS Attention are more concentrated, the local areas of the eyes are weakened, more attention is being paid to the movement of facial muscles. For instance, col. 4 (Happy) emphasizes the cheek area, while col. 2 (Fear), col. 3 (Disgust), and col. 5 (Sad) focus on the cheekbone area. Additionally, both col. 6 (Anger) and col. 5 (Sad) highlight the point between the eyebrows.
@@ CONCLUSION
This paper proposes a feature refining network named Quadruplet Cross Similarity (QCS) to mine richer intrinsic information for Facial Expression Recognition (FER). The QCS comprises a closed-loop symmetric network for stable joint training. It adaptively extracts discriminative features within the same class and synchronously separates interfering features between different classes. The cross module is based on Cross Similarity Attention (CSA), a novel global spatial attention designed for different image pairs, which directly assigns similarity attention between the input and output feature vectors of a branch. Extensive experiments on several public datasets indicate that our QCS effectively refines the features into a purer form and achieves state-of-the-art performance. Furthermore, effectively transferring the performance of the cross module back to the basic inference network while avoiding overfitting, and enhancing the stability of joint training, remains a crucial task.
@@ Acknowledgments
This work was supported by funding from the National Natural Science Foundation of China (U21B2035, 82372218).
Appendix
Failed Attempt for the Triplet Network
Figure 10: Challenges in constructing a stable training Triplet Network.
Unlike Triplet loss (Schroff, Kalenichenko, and Philbin 2015), our network employs CSA in conjunction with a supervised classifier to achieve the purpose of bringing similar classes closer and pushing different classes farther away at the fine-grained feature level. If the neg branch is not connected to the same classifiers ${Cls_{base}}$ and ${Cls_{cross}}$ as the anchor and pos branches, gradient conflicts will arise among different branches during training. Due to the different inputs for the cross module S and the cross module D, they are unable to share weights. If the neg branch is frozen, as illustrated by the blue dashed arrow in Fig. 10, then the cross module D cannot be effectively learned. Despite conducting numerous experiments, we were unable to train an effective model.
Concat and Split on Spatial Dimension
[…]
Training Improvement and Results Details
As can be seen from Fig. 12(a), the distribution of samples across different categories within each dataset is highly unbalanced, and there are also significant variations in the proportion of training and testing samples among different datasets. Taking into account the potential overfitting issue in the cross module of the framework, we adopted different parameter settings and sampling strategies when training on different datasets. Fig. 12(b) displays the accuracy confusion matrices for each test set, where the accuracy of each category correlates with the number of samples in the training set.
(a) sample distributions
(b) confusion matrices
Figure 12: Details of sample distribution and confusion matrices of QCS on each dataset.
RAF-DB
The ratio of training to testing images for each category is approximately 4:1. Since the test set and the training set exhibit similar unbalanced distributions, the relatively lower accuracy rates for the categories of Disgust and Fear do not significantly affect the overall accuracy. For positive samples, random sampling is employed. After determining the positive samples in QCS, negative samples are randomly sampled from all samples of the other categories. The learning rate is initialized as 9e-6 for DCS and 4e-6 for QCS. When fine-tuning QCS on the pre-trained model QCS-Affect-8, the learning rate is initialized to 3e-6.
FERPlus
The ratio of training to testing images for each category is approximately 8:1. Although, like the RAF-DB dataset, the test set and the training set exhibit similar unbalanced distributions, the extreme scarcity of samples in the Disgust and Contempt categories may exacerbate the overfitting issue specifically on these categories. The test results from the confusion matrix also indicate that the model’s accuracy is very low for the Disgust, Fear, and Contempt categories.
For positive samples, random sampling is employed. After determining the positive samples in QCS, negative samples are randomly sampled from all samples of the other categories. To mitigate overfitting, we intensified our sample augmentation efforts by performing random crops of 224x224 within images of 234x234, applying random rotations of up to 10 degrees, and randomly jittering brightness, contrast, saturation, and hue. Additionally, we set the dropout rate of ${{ViT}_{cross}}$ to 0.2 to stabilize the training results for DCS and QCS. The learning rate is initialized as 8e-6 and 3.5e-6 for QCS. When fine-tuning QCS on the pre-trained model QCS-Affect-8, the learning rate is initialized to 2.5e-6.
AffectNet-7/8
Given that the amount of training data has increased significantly compared to RAF-DB and FERPlus, we train model on the subset (only sample 1/N of the total samples) in one epoch, where N is 8 for DCS and 12 for QCS. The training will run for a total of 80 epochs.
Balance-sampling(Imb 2019) is used since the training set is unbalanced, while the validation set is balanced. Positive samples are selected using Balance-sampling(Imb 2019). After determining the categories of positive sample in QCS, negative sample categories are randomly chosen with equal probability from the remaining categories, and then negative samples are randomly sampled within the selected category to achieve Balance-sampling for negative samples.
Over-sampling minority classes in such a long-tailed distributed training set is prone to overfitting. To mitigate overfitting, we incorporate samples of the Fear, Disgust, and Contempt classes from RAF-DB and FERPlus for training. The dropout rate of ${{ViT}_{cross}}$ is set to 0.4 for DCS. We intensified our sample augmentation efforts by performing random crops of 224x224 within images of 236x236, applying random rotations of up to 12 degrees, and randomly jittering brightness, contrast, saturation, and hue. The learning rate is initialized as 9e-6 for DCS and 4e-6 for QCS.
Comparison without landmark
Methods Backbone RAF-DB FERPlus AffectNet-7
POSTER(w/o lm) IR50 90.51 - 64.95
POSTER++(w/o lm) IR50 91.62 - -
S2D(w/o lm) ViT-B/16 91.07 90.18 66.09
QCS(Ours) IR50 92.50 91.41 67.94
Table 4: Performance comparison (%), (w/o lm) means without landmark.
Figure 13: QCS Attention visualization for anchor. The pos samples all use the same image, while the neg samples each use an image from different categories.
Table 4 compares the performance of our QCS to POSTER, POSTER++ and S2D, all without incorporating additional landmark information. Our QCS has demonstrated significant performance improvements. Compared with POSTER++, our method shows no increase in inference time when all interactive modules are removed and only a single branch is retained.
Attention Visualization
Fig. 13 displays the extraction of distinctive features from two images belonging to the same category (Surprise), while simultaneously isolating redundant features across samples from six other distinct categories to highlight the similarities and differences in attention. Almost all images exhibit enhanced areas between the nose and the corners of the mouth, likely due to significant variations in muscle movement within this region across different expressions. On the other hand, the eye region in column 1 (Fear) resembles that of Surprise and is notably weakened.